# On a fast consistent selection of nested models with possibly unnormalised probability densities

### v2

@@ Abstract
Models with unnormalized probability density functions are ubiquitous in statistics, artificial intelligence and many other fields. However, they face significant challenges in model selection if the normalizing constants are intractable. Existing methods to address this issue often incur high computational costs, either due to numerical approximations of normalizing constants or evaluation of bias corrections in information criteria. In this paper, we propose a novel and fast selection criterion, T-GIC, for nested models, allowing direct data sampling from a possibly unnormalized probability density function. T-GIC gives a consistent selection under mild regularity conditions and is computationally efficient, benefiting from a multiplying factor that depends only on the sample size and the model complexity. Extensive simulation studies and real-data applications demonstrate the efficacy of T-GIC in the selection of nested models with unnormalized probability densities.
Keywords: unnormalized probability densities, gradient-based information criterion, consistent model selection, computational efficiency, nested models
@@ 1 Introduction
Models with unnormalized probability density functions are ubiquitous. In statistics, artificial intelligence, statistical mechanics and many other fields, often we only want or are able to stipulate the general shape of the models’ distributions without requiring their underlying probability density functions (PDFs) to integrate to unity, hence the notion of unnormalized PDFs. In such cases, the normalizing constants are either extremely difficult or impossible to compute directly. They arise in various circumstances, e.g. non-conjugacy in Bayesian posteriors (Andrade and Rathie, 2017), partition functions in statistical mechanics (Frigg and Werndl, 2024), directional distributions of data on a sphere (Pewsey and García-Portugués, 2021), Ising model in spatial statistics (Friel, 2013), and distributions with high-dimensional latent variables (Murray and Salakhutdinov, 2008). In short, unnormalized PDFs pose a significant challenge in likelihood-based model comparison. Addressing the challenge, we focus on the selection of nested models with possibly unnormalized PDFs.
@@
Model selection involves choosing the best statistical model from several candidates based on the observed data, with each candidate having possibly a different number of parameters (Rao et al., 2001). Among the forerunners and the most prominent members of this vast industry, Akaike (1974) proposed the Akaike Information Criterion (AIC) based on the Kullback-Leibler divergence that is underpinned by the Shannon entropy. Here, the model with the minimum AIC is selected as the best model. As the other, Schwarz (1978) developed the Bayesian Information Criterion (BIC) via a Laplace approximation of the Bayes factor. Here, the model with the minimum BIC value is selected. Unfortunately, none of them is applicable for the selection of models with unnormalized PDFs.
@@
At present, there are essentially two main approaches to overcome the challenge. One approach is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo. See, for example, (Baker, 2022) and (MacKay, 2003; Congdon, 2006). While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings. Another approach is to avoid the calculation of the normalizing constant altogether. For this, the score matching method (Hyvärinen, 2005) provides significant potential through efficient computation, as it does not depend on the normalizing constant. This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution. In this approach, consistent selection is only proved for non-nested models (Shao et al., 2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear. Our approach is different. We allow the data to be sampled directly from an unnormalized PDF.
@@
Recently, Matsuda et al. (2021) proposed an information criterion for selection of models with unnormalized PDFs estimated via noise contrastive estimation (NCE) or score matching. Slightly later, Cheng and Tong (2024) introduced a new entropy, the G-entropy, that includes an information criterion for model selection via a negative Hyvärinen score. In the above information criteria, bias correction is involved and the estimation of the bias term can be computationally intensive. For example, with parameter dimension $k$ and sample size $n$, matrix calculations for the bias in Cheng and Tong (2024) run in $O(nk^{2}+k^{3})$ time. Similar order applies to Matsuda et al. (2021)
@@
In this paper, we propose a fast information criterion for consistent selection of nested models with unnormalized PDFs. We name the criterion T-GIC. In it, we introduce a multiplying factor $C(n,k)$, which depends only on the sample size $n$ and the order $k$ of the candidate model $M_{k}$, resulting in a drastic reduction of computation to $O(1)$. We show that T-GIC achieves consistent model selection for strictly nested models under mild regularity conditions. Further, we demonstrate the efficacy of T-GIC through simulation on AR and polynomial regression models with unnormalized PDFs as well as a model with a bivariate von Mises PDF with bounded support. Finally, we apply the T-GIC to real data from diverse domains, including finance, automotive engineering, and wind direction analysis.
This paper is organized as follows. In Section 2, we provide a brief review of F-divergence and associated notions. In Section 3, we introduce the T-GIC and prove its consistency for model selection under regularity assumptions. In Section 4, we conduct simulations to demonstrate the efficacy of T-GIC. In Section 5, we apply T-GIC to various real-world data, offering some insights. In Section 6, we conclude with a discussion of our findings and potential directions for future research. Proofs are provided in Appendix A.
@@ 2 A New Information Criterion
It is well known that the Kullback-Leibler divergence is underpinned by the Shannon entropy. Recently, Cheng and Tong (2024) have developed a new entropy that parallels Shannon’s. It is called the G-entropy (sometimes shortened to Gentropy). They have shown how it underpins the Fisher divergence and leads to the GIC for model selection.
2.1 Fisher Divergence and Gentropy
Let $p(x)$ be a PDF on $\mathbb{R}^{d}$ under the following assumptions:
@@ 2.1 Fisher Divergence and Gentropy
$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0,$
@@
Let $q(x)$ be a PDF on $\mathbb{R}^{d}$ satisfying Assumptions 1 and 2. The Fisher divergence between PDF $p$ and PDF $q$ is defined as
$D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}.$ (2.1)
The objective function $W$ of PDF $q$ at point $x$ is a negative Hyvärinen score (Hyvärinen, 2005), defined by
$W(x,q)=-||\nabla_{x}\log q(x)||^{2}~{}-~{}2\Delta_{x}\log q(x),$ (2.2)
where $\Delta$ denotes the Laplacian, i.e. $\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$.
Next, within their new entropy structure, Cheng and Tong (2024) defined, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
$H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2},$ (2.3)
$H_{G}(p,q)=E_{p}[W(x,q)].$ (2.4)
Integration by parts yields the following relationship
$D_{F}(p||q)=H_{G}(p)-H_{G}(p,q).$
Denote $x$ and $y$ as two random vectors with joint PDF $p(x,y)$. Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$. Tables 1 and 2 summarize a comparison between the Shannon entropy and the G-entropy (in terms of definitions and properties). The tables are based on the results developed by Cheng and Tong (2024).
Table 1: Comparison between Shannon entropy and G-entropy: definitions.
Definitions Shannon entropy G-entropy
Objective function $Loss(p(x))=-\log p(x)$ $W(x,p)=-||\nabla_{x}\log p(x)||^{2}~{}-~{}2\Delta_{x}\log p(x)$
Entropy $H(x)=H(p)=E_{p}[-\log p(x)]$ $H_{G}(x)=H_{G}(p)=E_{p}[W(x,p)]$
Cross entropy $H(p,q)=E_{p}[-\log q(x)]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
Joint entropy $H(x,y)=H(p_{(x,y)})$ $H_{G}(x,y)=H_{G}(p_{(x,y)})$
Conditional entropy $H(y|x)=E_{p_{(x,y)}}[Loss(\frac{p(x,y)}{p(x)})]$ $H_{G}(y|x)=E_{p(x,y)}[W((x,y),\frac{p(x,y)}{p(x)})]$
Mutual information $I(x,y)=H(x)+H(y)-H(x,y)$ $I_{G}(x,y)=H_{G}(x,y)-H_{G}(x)-H_{G}(y)$
Information matrix $IM_{\theta}=E_{p}[[\nabla_{\theta}\log p(x)][\nabla_{\theta}\log p(x)]^{\tau}]$ $GIM_{x}=E_{p}[[\nabla_{x}\log p(x)][\nabla_{x}\log p(x)]^{\tau}]$
Divergence $D_{KL}(p||q)=E_{p}[Loss(\frac{q(x)}{p(x)})]$ $=H(p,q)-H(p)$ $D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}$ $=H_{G}(p)-H_{G}(p,q)$
Table 2: Comparison between Shannon entropy and G-entropy: properties.
Shannon entropy G-entropy
1 $H(p,p)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(p,p)=H_{G}(p)=E_{p}[W(x,p)]$
2 $H(p)\leq H(p,q)$ $H_{G}(p)\geq H_{G}(p,q)$
3 $\min_{q}\{H(p,q)\}=H(p)$ $\max_{q}\{H_{G}(p,q)\}=H_{G}(p)$
4 $H(x,y)\leq H(x)+H(y)$, with equality iff $x$ and $y$ are independent. $H_{G}(x,y)\geq H_{G}(x)+H_{G}(y)$, with equality iff $x$ and $y$ are independent.
5 $H(y|x)=H(x,y)-H(x)$ $H_{G}(y|x)=H_{G}(x,y)-H_{G}(x)$
@@ 2.2 GIC Estimate
By equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$. When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
@@
$\theta^{*}=\arg\min_{\theta\in\Theta}D_{F}(p_{x}||p_{M(\theta)})=\arg\max_{\theta\in\Theta}H_{G}(p_{x},p_{M(\theta)}).$ (2.5)
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. An unbiased estimate of $H_{G}(p_{x},p_{M(\theta)})$ is given by
@@ 2.3 GIC for Model Selection
Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
@@
3 A New Model Selection Criterion: T-GIC
In this section, we propose a fast model selection criterion, T-GIC, and show the consistency for a finite sequence of strictly nested models under mild regularity conditions.
@@ 3.1 A Fast Model Selection
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. Suppose we have a collection of candidate parametric models as described in Section 2.3. The $T\text{-}GIC(k)$ of model $M_{k}$ is defined as
$T\text{-}GIC(k)=C(n,k)\times GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
where $C(n,k)$ is a constant depending only on $n$ and $k$, $GIC(M_{k}(\hat{\theta}_{k}))$ is defined by equation (2.6), and $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$ for model $M_{k}$.
We propose to select the model that maximizes $T\text{-}GIC(k)$. Compared to $GIC_{c}$ in equation (2.8), $T\text{-}GIC$ significantly reduces the computational costs by introducing a factor $C(n,k)$ to bypass the bias correction calculation. Note that estimating the bias, B, based on a sample of size $n$ involves a calculation $O(nh_{k}^{2}+h_{k}^{3})$, which becomes computationally expensive when $h_{k}$ or $n$ is large. In contrast, using the multiplying factor $C(n,k)$ reduces the computational cost to only $O(1)$. The following section shows the consistency of the proposed model selection criterion T-GIC.
3.2 Consistency of T-GIC
We first introduce several assumptions, which have been used by Song et al. (2020) and Cheng and Tong (2024).
@@ 3.2 Consistency of T-GIC
Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm. Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
$||\nabla_{x}^{2}\log p_{M(\theta_{1})}(x)-\nabla_{x}^{2}\log p_{M(\theta_{2})}(x)||_{F}\leq L_{1}(x)||\theta_{1}-\theta_{2}||_{2},$
@@
$||[\nabla_{x}\log p_{M(\theta_{1})}(x)][\nabla_{x}\log p_{M(\theta_{1})}(x)]^{T}-[\nabla_{x}\log p_{M(\theta_{2})}(x)][\nabla_{x}\log p_{M(\theta_{2})}(x)]^{T}||$
$\leq L_{2}(x)||\theta_{1}-\theta_{2}||_{2}.$
@@
$||\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq M_{i,j}(x)||\theta_{1}-\theta_{2}||_{2},$
@@
$||\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq N_{i,j}(x)||\theta_{1}-\theta_{2}||_{2}.$
@@
Note that Assumptions 4 and 6 are standard conditions for proving the consistency of the maximum likelihood estimation (MLE). Assumption 5 is also used by Hyvärinen (2005). Assumption 7 defines Lipschitz continuity, while Assumption 8 describes Lipschitz smoothness for second derivatives. Based on these assumptions, Proposition 6 in Cheng and Tong (2024) shows the asymptotic normality of the MGICE, as stated in Lemma 1.
@@
Objective function $\log[p_{M(\theta)}(x_{i})]$ $W(x_{i},p_{M(\theta)})=-||\nabla_{x}\log p_{M(\theta)}(x_{i})||^{2}~{}-~{}2\Delta_{x}\log p_{M(\theta)}(x_{i})$
@@
Now, we establish the consistency of our T-GIC method.
@@
Let $\hat{k}=\arg\max_{1\leq k\leq K}T\text{-}GIC(k)$. Then $\hat{k}$ converges to $k_{0}$, in probability.
@@
where $p_{M(\theta)}(\cdot|x_{i-1},\cdots,x_{i-k})$ is the conditional pdf of $x_{i}$ given its lags 1 to $k$. It follows from Assumptions 1–3 that $\nabla_{\theta}CGIC(M(\theta))$, evaluated at the true parameter, is a martingale difference sequence. By appealing to some variant of the martingale central limit theorem (Hall and Heyde, 2014, Chapter 3), the proofs of Proposition 1 and Theorem 1 for the IID case can be easily extended to the case of stationary ergodic finite-order Markov processes, with MGICE and T-GIC modified accordingly.
Mimicking AIC and BIC, we set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively, where #$\#(M_{k})$ is the number of independently adjusted parameters in model $M_{k}$. From Theorem 1, the following corollaries follow immediately.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
However, similar to AIC, T-GIC1 may fail to provide consistent selection when $K>k_{0}$. In the following sections, we will compare the performance of T-GIC1 and T-GIC2 through simulations and applications.

### v3

# On a fast consistent selection of nested models with possibly unnormalized probability densities

@@ Abstract
Models with unnormalized probability density functions are ubiquitous in statistics, artificial intelligence and many other fields. However, they face significant challenges in model selection if the normalizing constants are intractable. Existing methods to address this issue often incur high computational costs, either due to numerical approximations of normalizing constants or evaluation of bias corrections in information criteria. In this paper, we propose a novel and fast selection criterion, MIC, for nested models of possibly dependent data, allowing direct data sampling from a possibly unnormalized probability density function. With a suitable multiplying factor depending only on the sample size and the model complexity, MIC gives a consistent selection under mild regularity conditions and is computationally efficient. Extensive simulation studies and real-data applications demonstrate the efficacy of MIC in the selection of nested models with unnormalized probability densities.
Keywords: Unnormalized probability densities, Gradient-based information criterion, Consistent model selection, Computational efficiency, Nested models, Markov dependent data
@@ 1 Introduction
Models with unnormalized probability density functions are ubiquitous. In statistics, artificial intelligence, statistical mechanics and many other fields, often we only want or are able to stipulate the general shape of the models’ distributions without requiring the underlying probability density functions (PDFs) to integrate to unity, hence the notion of unnormalized PDFs. Here, the normalizing constants are either difficult or impossible to compute explicitly. They arise in various circumstances, e.g. non-conjugacy in Bayesian posteriors (Andrade & Rathie 2017), partition functions in statistical mechanics (Frigg & Werndl 2024), directional distributions of data on a sphere (Pewsey & García-Portugués 2021), Ising model in spatial statistics (Friel 2013), and distributions with high-dimensional latent variables (Murray & Salakhutdinov 2008). In short, unnormalized PDFs pose a significant challenge in likelihood-based model comparison. Addressing it, we focus on the selection of nested models with possibly unnormalized PDFs.
@@
Model selection involves choosing the best statistical model from several candidates based on the observed data, with each candidate having possibly a different number of parameters (Rao et al. 2001). Among the forerunners, Akaike (1974) proposed the Akaike Information Criterion (AIC) based on the Kullback-Leibler divergence. As the other, Schwarz (1978) developed the Bayesian Information Criterion (BIC) via a Laplace approximation of the Bayes factor. Unfortunately, none of them is applicable for the selection of models with unnormalized PDFs.
@@
There are two main approaches to overcome the challenge. One is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo. See, for example, (Baker 2022) and (MacKay 2003, Congdon 2006). While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings. Another is to avoid the calculation of the normalizing constant altogether via the score matching method (Hyvärinen 2005) that provides significant potential through efficient computation without the normalizing constant. This method has been applied to Bayesian model selection (Dawid & Musio 2015, Shao et al. 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution. In this approach, consistent selection is only proved for non-nested models (Shao et al. 2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear.
@@
Direct sampling from an unnormalized PDF in model selection is challenging. For this, the following approaches are available. Recently, for independent data, Matsuda et al. (2021) proposed an information criterion for selection of models with unnormalized PDFs estimated via noise contrastive estimation (NCE) or score matching. Slightly later, for both independent and Markov dependent data, Cheng & Tong (2024) proposed an information criterion for model selection. In the above information criteria, bias correction is involved and the estimation of the bias term can be computationally intensive. For example, with parameter dimension $k$ and sample size $n$, matrix calculations for the bias in Cheng & Tong (2024) run in $\mathcal{O}(nk^{2}+k^{3})$ time. Similar order applies to Matsuda et al. (2021).
@@
In this paper, we propose a fast information criterion for consistent selection of nested models of possibly Markov dependent data with unnormalized PDFs. We name the criterion MIC. In it, we introduce a multiplying factor $C(n,k)$, which depends only on the sample size $n$ and the order $k$ of the candidate model $M_{k}$, resulting in a drastic reduction of computation to $\mathcal{O}(1)$. We show that MIC achieves consistent model selection for strictly nested models under mild regularity conditions. Further, we demonstrate the efficacy of MIC through simulation on AR and polynomial regression models with unnormalized PDFs as well as a model with a bivariate von Mises PDF with bounded support. Finally, we apply the MIC to real data from diverse domains, including finance, automotive engineering, and wind direction analysis.
[…]
@@ 2 A New Information Criterion
2.1 Fisher Divergence
Let $p(x)$ be a generic PDF on $\mathbb{R}^{d}$ under the following assumptions:
@@ 2.1 Fisher Divergence
[…]
@@
Consider the following objective function:
[…]
where $\Delta$ denotes the Laplacian, i.e. $\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$. For PDFs $p$ and $q$ under the above assumptions, the Fisher divergence satisfies the following equation:
$D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~-~\nabla_{x}\log q(x)||^{2}=E_{p}[W(x,p)]-E_{p}[W(x,q)].$ (2.2)
See, e.g., Cheng & Tong (2024). Note that $E_{p}[W(x,p)]$, interpreted as an entropy, is $trace(G_{p}),$ where $G_{p}$ is a matrix whose $(i,j)$-th element is $E_{p}[f_{i}(x)f_{j}(x)/p^{2}(x)],$ $f_{i}(x)=\partial p/\partial x_{i}.$ See (Cheng & Tong (to appear in 2025)), who introduced the matrix under the name covariate Fisher Information Matrix based on a Riemannian geometry approach.
@@ 2.2 GIC Estimate
Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$. When $p$ is the true data PDF and $q$ the model PDF, by equation (2.2), given PDF $p$ minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing $E_{p}[W(x,q)],$ inspiring the maximum GIC estimate (MGICE).
@@
$\theta^{*}=\arg\min_{\theta\in\Theta}D_{F}(p_{x}||p_{M(\theta)})=\arg\max_{\theta\in\Theta}E_{p}[W(x,p_{M(\theta)})].$ (2.3)
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. An unbiased estimate of $E_{p}[W(x,p_{M(\theta)})]$ is given by
@@ 2.3 GIC for Model Selection
Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ An unbiased $GIC_{c}$ for model selection is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
@@
3 A New Model Selection Criterion: MIC
In this section, we propose a fast model selection criterion, MIC, and show the consistency for a finite sequence of strictly nested models under mild regularity conditions.
@@ 3.1 A Fast Model Selection
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. Suppose we have a collection of candidate parametric models as described in Section 2.3. The $MIC(k)$ of model $M_{k}$ is defined as
$MIC(k)=C(n,k)\times GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
where $C(n,k)$ is a constant depending only on $n$ and $k$, $GIC(M_{k}(\hat{\theta}_{k}))$ is defined by equation (2.4), and $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$ for model $M_{k}$. We propose to select the model that maximizes $MIC(k)$. The following sections demonstrate the high computational efficiency and the selection consistency of the proposed criterion $MIC$.
3.2 Computational cost comparison
[…]
Table 1: Computational cost of penalty terms in model selection criteria.
Criteria GICc NCIC1 NCIC2 MIC
[…]
3.3 Consistency of MIC
The following assumptions are from Song et al. (2020) and Cheng & Tong (2024).
@@ 3.3 Consistency of MIC
Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm. Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
[…]
@@
$\nabla_{x}\log p_{M(\theta_{1})}(x)][\nabla_{x}\log p_{M(\theta_{1})}(x)]^{T}-[\nabla_{x}\log p_{M(\theta_{2})}(x)][\nabla_{x}\log p_{M(\theta_{2})}(x)]^{T}||$
[…]
@@
[…]
@@
[…]
@@
Note that Assumptions 4 and 6 are standard conditions for proving the consistency of the maximum likelihood estimation (MLE). Assumption 5 is also used by Hyvärinen (2005). Assumption 7 defines Lipschitz continuity, while Assumption 8 describes Lipschitz smoothness for second derivatives. Based on these assumptions, Proposition 6 in Cheng & Tong (2024) shows the asymptotic normality of the MGICE, as stated in Lemma 1.
@@
[…]
@@
Now, we establish the consistency of our MIC method.
@@
Let $\hat{k}=\arg\max_{1\leq k\leq K}MIC(k)$. Then $\hat{k}$ converges to $k_{0}$, in probability.
@@
where $p_{M(\theta)}(\cdot|x_{i-1},\cdots,x_{i-k})$ is the conditional pdf of $x_{i}$ given its lags 1 to $k$. It follows from Assumptions 1–3 that $\nabla_{\theta}CGIC(M(\theta))$, evaluated at the true parameter, is a martingale difference sequence. By appealing to some variant of the martingale central limit theorem (Hall & Heyde 2014, Chapter 3), the proofs of Proposition 1 and Theorem 1 for the IID case can be easily extended to the case of stationary ergodic finite-order Markov processes, with MGICE and MIC modified accordingly.
Let us consider the following choices of $C(n,k)$: (1) $\exp\{-2\#(M_{k})/n\}$ and (2) $n^{-\#(M_{k})/n}$,
where $(M_{k})$ is the number of independently adjusted parameters in model $M_{k}$. From Theorem 1, the following corollaries follow immediately.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. If $K=k_{0}$, under Assumptions 1-8, then $MIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
Note that MIC1 may fail to provide consistent selection when $K>k_{0}$.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Under Assumptions 1-8, then $MIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
In the following sections, we will compare the performance of MIC1 and MIC2 through simulations and applications.